# Bilateral facial reduction: qualification-free subdifferential calculus and exact duality in convex analysis

### v1

@@ 2 Prior works
One theoretical insight that was critical to developing our theory was the lexicographical characterization of facial subspaces from [8]. This work shows that every face of a convex set can be obtained by iteratively exposing faces. It is a sequel of a preceding result [9], which shows a separation theorem for convex sets. Though the lexicographical framework was not flexible enough for our needs, prompting us to develop an independent theory, we did draw inspiration from the results and the proofs. In this paper, we recover this lexicographical representation, though in different terms, primarily in Lemma 6.5. Our proof is direct, and does not make use of a separation theorem.
@@ 3 Preliminaries
For a set $S\subseteq\mathbb{R}^{n}$, denote by $\operatorname{conv}(S)$ its convex hull. For $y_{1},y_{2}\in\mathbb{R}^{n}$, let $y_{1},y_{2}]:=\operatorname{conv}\{y_{1},y_{2}\}$, and $y_{1},y_{2}[:=\operatorname{conv}\{y_{1},y_{2}\}\setminus\{y_{1},y_{2}\}$.
Intuitively the notion of a face of a convex set $C$ is a subset of $C$ for which there is no line segment in $C$ that “goes through” $F$. The following definitions can be found inline in [16, p. 16] and [16, p.75] respectively.
@@
We also recall some objects from convex analysis. Let $\bar{\mathbb{R}}=\mathbb{R}\cap\{-\infty,\infty\}$. For a convex function $f:\mathbb{R}^{n}\to\bar{\mathbb{R}}$, the subdifferential is
$\partial f(x):=\{v\in\mathbb{R}^{n}\mid\forall y\in\mathbb{R}^{n},\langle v,y-x\rangle\leq f(y)\},$
and the convex conjugate is $f^{*}(y):=\sup_{x\in\mathbb{R}^{n}}\langle y,x\rangle-f(x)$.
4 The joint facial subspace
@@
Definition 4.1 (The joint facial subspace).
The joint facial subspace of two convex sets $C,D\subseteq\mathbb{R}^{n}$ is
@@
We also define $T_{a}(C,D):=T(C,D)+C\cap D$, the joint affine facial subspace.
The Minkowski sum $(+C\cap D)$ in the definition of $T_{a}$ is in fact a simple translation. This follows from the next corollary.
@@
Let $C,D\in\mathbb{R}^{n}$ be convex. Then
@@
This result follows from Lemma 6.10; we defer the proof of Corollary 4.1 to the end of Section 6.2.
From Corollary 4.1, we see that picking any $x\in C\cap D$, we have $T_{a}=T+x$, independently of the choice of $x$. Corollary 4.1 also implies that $C\cap D\subseteq T_{a}$. The affine subspace $T_{a}$ is therefore the translation of $T$ that contains $C\cap D$.
In this section we provide two characterizations of the joint facial subspace. The next result shows the relation of the joint facial subspace to the individual convex sets, and builds towards the first characterization.
Proposition 4.1 (The joint facial subspace reveals faces).
@@
For the first characterization we require a subspace that is related to $F_{C}(C\cap D)$. This subspace, the generated facial subspace, is central to our theory, as is made clear in Section 6. There, we will re-introduce this object in a motivated way, in Definition 6.5.
Definition 4.2 (Preliminary definition of the generated facial subspace).
Given a convex set $C\subseteq\mathbb{R}^{n}$ with $S\subseteq C$ non-empty, define the facial subspace of $C$ generated by $S\subseteq C$ to be
@@
The next theorem is our first characterization of the joint facial subspace.
@@
The joint facial subspace of $C,D\subseteq\mathbb{R}^{n}$ is
@@
Definition 4.1 and Theorem 4.1 both specify the joint facial subspace via analytical formulas. Our second characterization instead specifies the joint facial subspace as the result of a process with a number of steps bounded by the ambient dimension, which is more amenable to computation. This result, Theorem 7.1, is best stated in terms of nested normals a concept that we introduce in Definition 6.3, and motivate throughout Section 6. For this reason, we defer it to Section 7.3. For the moment, we state a direct corollary which does not require this object.
@@
Now that the joint facial subspace has been thoroughly identified, we move on to describing its core property. Recall that for two convex functions $f,g:\mathbb{R}^{n}\to\bar{\mathbb{R}}$, a common qualification condition is that for convex sets $C:=\operatorname{dom}(f),D:=\operatorname{dom}(g)$,
@@
It was shown in [14, theorem 11.3] that this notion is equivalent to a lack of proper separation of the domains. We define the notions of separation and proper separation below.
@@
Two convex sets $C,D\subseteq\mathbb{R}^{n}$ are said to be separated if there exists some $v\in\mathbb{R}^{n}\setminus\{0\}$ such that
$\forall c\in C,d\in D,v\cdot c\leq v\cdot d.$ (2)
@@
Sets $C,D\in\mathbb{R}^{n}$ are said to be properly separated when there exists some $v\in\mathbb{R}^{n}$ such that Equation 2 holds, and such that there is some pair $(c,d)\in C\times D$ for which the inequality in Equation 2 is strict.
Proper separation excludes the case where the two sets are merely separated by virtue of being both contained in some affine hyperplane.
We will show that inside the joint facial subspace, we always have a “lack of separation” of the convex sets. Since this regularity condition is stronger than a “lack of proper separation”, it is stronger than Equation 1.
Theorem 4.2 (Qualification conditions always holds in the joint facial subspace).
For two convex sets $C,D\subseteq\mathbb{R}^{n}$, $C\cap T_{a}(C,D)$ and $D\cap T_{a}(C,D)$ are not separated.
The proof of Theorem 4.2 is the subject of Section 7.1. The next corollary follows immediately from Theorem 4.2.
Corollary 4.3 (rint qualification holds in joint facial subspace).
For convex sets $C,D\subseteq\mathbb{R}^{n}$,
@@
With Corollary 4.3 we can use standard results of convex analysis from [14] by first taking intersections with $T_{a}$. We do this in the proof of each of the results in Section 5. Also note that if we only had access to weaker results than those available in [14], results that would have the stronger assumption of “no separation of the sets” instead of “intersection of relative interiors is non-empty”, the theory in this paper would still hold, and indeed we would recover the stronger results with the weaker assumption, that the intersection of relative interiors is non-empty.
@@ 5 Main results
In this section we show five generalizations of important results of convex analysis: additivity of subdifferentials, normal cone of the intersection of two convex sets, attainment of the infimal convolution in the conjugate of a sum of two functions, the chain rule of the subdifferential, and the Fenchel-Rockafellar dual. All proofs are deferred to Section 7, because they require the machinery introduced in Section 6.
@@
In all results that follow, the classical analogue is recovered when qualification conditions hold. Indeed, given two convex functions $f,g$ such that their relative interiors intersect, we have $\operatorname{dom}(f),\operatorname{dom}(g)\subseteq T_{a}(\operatorname{dom}(f),\operatorname{dom}(g))$, and therefore $f+\iota_{T_{a}}=f$ and $g+\iota_{T_{a}}=g$.
For a proper convex function $f$, the singular subdifferential for some $x\in\mathbb{R}^{n}$ is defined as $\partial^{\infty}f(x):=N_{\operatorname{dom}(f)}(x)$ when $x\in\operatorname{dom}(f)$, and is empty when $x\notin\operatorname{dom}(f)$ (see, e.g. [10, 15]). Singular subdifferentials capture the notion of an infinite slope in a particular direction.
Theorem 5.1 (Qualification-free subdifferential additivity).
For $f,g:\mathbb{R}^{n}\to\bar{\mathbb{R}}$ proper convex, $x\in\operatorname{dom}(f)\cap\operatorname{dom}(g)$, and $T_{a}=T_{a}(\operatorname{dom}(f),\operatorname{dom}(g))$,
$\partial(f+g)(x)=\partial(f+\iota_{T_{a}})(x)+\partial(g+\iota_{T_{a}})(x),$
$\partial^{\infty}(f+g)(x)=\partial^{\infty}(f+\iota_{T_{a}})(x)+\partial^{\infty}(g+\iota_{T_{a}})(x).$
Proof of Theorem 5.1.
By Corollary 4.3 applied to $\operatorname{dom}(f)$ and $\operatorname{dom}(g)$,
$\operatorname{relint}\operatorname{dom}(f+\iota_{T_{a}})\cap\operatorname{relint}\operatorname{dom}(g+\iota_{T_{a}})$ (3)
$=$ $\operatorname{relint}(\operatorname{dom}(f)\cap T_{a})\cap\operatorname{relint}(\operatorname{dom}(g)\cap T_{a})\neq\emptyset.$ (4)
Therefore, a qualification condition is satisfied for the functions $f+\iota_{T_{a}}$ and $g+\iota_{T_{a}}$, and the first part of the result follows by application of [14, Theorem 23.8].
The second part of the result follows from applying [14, Corollary 23.8.1] to the convex sets $\operatorname{dom}(f)\cap T_{a}$ and $\operatorname{dom}(g)\cap T_{a}$, by which
$N_{(\operatorname{dom}(f)\cap T_{a})\cap(\operatorname{dom}(g)\cap T_{a})}(x)$ $=N_{\operatorname{dom}(f)\cap T_{a}}(x)+N_{\operatorname{dom}(g)\cap T_{a}}(x)$
$=N_{\operatorname{dom}(f+\iota_{T_{a}})}(x)+N_{\operatorname{dom}(g+\iota_{T_{a}})}(x).$
And since $N_{(\operatorname{dom}(f)\cap T_{a})\cap(\operatorname{dom}(g)\cap T_{a})}(x)=N_{\operatorname{dom}(f)\cap\operatorname{dom}(g)}(x)=N_{\operatorname{dom}(f+g)}(x)$, the second part of the result follows. ∎
Theorem 5.2 (Normal cone of the intersection of convex sets).
For $C,D\subseteq\mathbb{R}^{n}$ convex, and $x\subseteq C\cap D$, for $T_{a}=T_{a}(C,D)$,
@@
Proof of Theorem 5.2.
With Corollary 4.3 we can use [14, Corollary 23.8.1], from which we find that
@@
$(f\,\square g)(z):=\inf_{x,y\in\mathbb{R}^{n}:x+y=z}f(x)+g(y).$
In the following results, we denote by $f^{*}$ the convex conjugate of $f$.
Theorem 5.3 (Attained infimal convolution in the conjugate of a sum).
For two proper convex functions $f,g:\mathbb{R}^{n}\to\bar{\mathbb{R}}$, and $T_{a}=T_{a}(\operatorname{dom}(f),\operatorname{dom}(g))$,
$(f+g)^{*}=(f+\iota_{T_{a}})^{*}\square(g+\iota_{T_{a}})^{*},$
where the infimum in the infimal convolution is attained.
Proof of Theorem 5.3.
When $\operatorname{dom}(f)\cap\operatorname{dom}(g)=\emptyset$, $T_{a}=\emptyset$ and the equality holds trivially. Otherwise, by Corollary 4.3, $\operatorname{relint}(\operatorname{dom}(f+\iota_{T_{a}}))\cap\operatorname{relint}(\operatorname{dom}(g+\iota_{T_{a}}))\neq\emptyset$, and the result follows from [14, Theorem 20.1]. ∎
Theorem 5.4 (Generalized subdifferential chain rule).
Let $g\in\mathbb{R}^{m}\to\bar{\mathbb{R}}$ be a proper convex function, and $A\in\mathbb{R}^{m\times n}$ is a matrix. Then for $T=T(\operatorname{dom}(g),\operatorname{\mathrm{range}}(A))$,
$\partial(g)(Ax)=A^{T}\partial(g+\iota_{T})(Ax).$
Proof of Theorem 5.4.
@@
If $\operatorname{dom}(g)\cap\operatorname{\mathrm{range}}(A)=\emptyset,$ then the result holds trivially because $g(Ax)\equiv\infty$, and $(g+\iota_{T_{a}})(Ax)\equiv\infty$ as well. Otherwise, by Corollary 4.3 $\operatorname{relint}\operatorname{dom}(g+\iota_{T_{a}})\cap\operatorname{\mathrm{range}}(A)\neq\emptyset$. Additionally, note that the only face of $\operatorname{\mathrm{range}}(A)$ is itself, whence by Proposition 4.1 $\operatorname{\mathrm{range}}(A)\subseteq T_{a}$. Since $0\in\operatorname{\mathrm{range}}(A)$, this in fact implies that $T_{a}=T$. The statement then follows by applying [14, Theorem 23.9] to the composition $g+\iota_{T}]\circ A$. ∎
In the following result, for a matrix $A\in\mathbb{R}^{m\times n}$ and a set $S\subseteq\mathbb{R}^{m}$, we denote by $A^{-1}(S)$ the preimage of $S$ under the linear map associated with the matrix $A$, making no assumption on the invertibility of the matrix $A$.
Theorem 5.5 (Exact Fenchel-Rockafellar dual).
For $f,g:\mathbb{R}^{n}\to\bar{\mathbb{R}}$ proper, lower semi-continuous convex functions, and let $T_{a}:=T_{a}(A\operatorname{dom}(f),\operatorname{dom}(g))$. Then
$\inf_{x\in\mathbb{R}^{n}}f(x)+g(Ax)=-\inf_{y\in\mathbb{R}^{m}}[(f+\iota_{A^{-1}(T_{a})})^{*}(A^{T}y)+(g+\iota_{T_{a}})^{*}(-y)].$

### v3

@@ 2.3 Theory of convex faces
A geometric insight that was critical to the development of our theory was the lexicographical characterization of faces [martinez-legazLexicographicalSeparationRn1987, martinez-legaz_LEXICOGRAPHICALCHARACTERIZATIONFACES_]. [martinez-legazLexicographicalSeparationRn1987] derive a separation theorem from which the characterization follows. It turned out that the lexicographical framework was not flexible enough for our needs, prompting us to develop an analogous yet independent and self-contained theory. In this paper, we recover the lexicographical characterization of convex faces in Theorem 6.4, though in different terms. Our proof is direct, and does not make use of a separation theorem, though some separation ideas are implicit in the definition of supporting subspaces (Definition 6.1).
@@ 3 Preliminaries
For a set $S\subseteq\mathbb{R}^{n}$, denote by $\operatorname{conv}(S)$ its convex hull (smallest containing convex set). For $y_{1},y_{2}\in\mathbb{R}^{n}$, let $y_{1},y_{2}]:=\operatorname{conv}\{y_{1},y_{2}\}$, and $y_{1},y_{2}[:=\operatorname{conv}\{y_{1},y_{2}\}\setminus\{y_{1},y_{2}\}$. We also denote $\operatorname{cone}(S)$ for the conical hull of a set; the smallest cone containing $S$.
Intuitively the notion of a face of a convex set $C$ is a subset of $C$ for which there is no line segment in $C$ that “goes through” $F$. The following definitions can be found inline in [schneider_ConvexBodiesBrunn_2013, p. 16] and [schneider_ConvexBodiesBrunn_2013, p.75] respectively.
@@
For $x\in\mathbb{R}^{n}$ and a convex set $C\subseteq\mathbb{R}^{n}$, the normal cone of $C$ at $x$ is
$N_{C}(x)=\{y\mid\forall c\in C,\langle y,c-x\rangle\leq 0\}$
if $x\in C$, and $N_{C}(x)=\emptyset$ otherwise.
We also recall some objects from convex analysis. Let $\bar{\mathbb{R}}=\mathbb{R}\cup\{-\infty,\infty\}$. For a convex function $f:\mathbb{R}^{n}\to\bar{\mathbb{R}}$, the subdifferential is
$\partial f(x):=\{v\in\mathbb{R}^{n}\mid\forall y\in\mathbb{R}^{n},f(x)+\langle v,y-x\rangle\leq f(y)\},$
with $\partial f(x)=\emptyset$ when $x\notin\operatorname{dom}(f)$. The convex conjugate is
$f^{*}(y):=\sup_{x\in\mathbb{R}^{n}}\langle y,x\rangle-f(x).$
4 The joint supporting subspace
Figure 1: In each figure, the green subspace is the joint supporting subspace for the red and blue convex sets.
@@ 4 The joint supporting subspace
Definition 4.1 (The joint supporting subspace).
The joint supporting subspace of two convex sets $C,D\subseteq\mathbb{R}^{n}$ is
@@
We also define its affine counterpart: $T_{a}(C,D):=T(C,D)+C\cap D$.
The Minkowski sum $(+C\cap D)$ in the definition of $T_{a}$ is in fact an affine translation. This follows from the next corollary.
@@
Let $C,D\subseteq\mathbb{R}^{n}$ be convex. Then
@@
All proofs for results presented in this section are deferred to Section 7, because they require the machinery introduced in Section 6. Corollary 4.1 follows from Theorem 6.6; for details, the proof can be found at the end of Section 6.2.
From Corollary 4.1, we see that picking any $x\in C\cap D$, we have $T_{a}=T+x$, independently of the choice of $x$. Corollary 4.1 also implies that $C\cap D\subseteq T_{a}$, showing that $T_{a}$ is the unique translation of $T$ that contains $C\cap D$.
In this section we provide two main characterizations of the joint supporting subspace. The next result builds towards the first characterization, describing the relation of the joint supporting subspace with the individual convex sets.
Lemma 4.1 (The joint supporting subspace reveals faces).
@@
Next we introduce the generated supporting subspace, which is central to our theory. This is made clear in Section 6, where we re-introduce this object in motivated fashion in Definition 6.5.
Definition 4.2 (Preliminary definition of the generated supporting subspace).
Given a convex set $C\subseteq\mathbb{R}^{n}$ with $S\subseteq C$ non-empty, define the supporting subspace of $C$ generated by $S\subseteq C$ to be
@@
The next theorem is our first characterization of the joint supporting subspace.
@@
The joint supporting subspace of $C,D\subseteq\mathbb{R}^{n}$ is
@@
We provide two other characterizations that follow from Theorem 4.1.
Corollary 4.2 (Joint supporting subspace as hull of faces).
For convex sets $C,D\subseteq\mathbb{R}^{n}$,
$T_{a}(C,D)=\operatorname{aff}(F_{C}(C\cap D)\cup F_{D}(C\cap D)).$
Corollary 4.3 (Joint supporting subspace as difference of faces).
For convex sets $C,D\subseteq\mathbb{R}^{n}$,
$T(C,D)=\operatorname{\mathrm{span}}(F_{C}(C\cap D)-F_{D}(C\cap D)).$
So far, we have characterized the joint supporting subspace with various analytical formulas. Our next characterization instead specifies how to obtain the joint supporting subspace constructively, from an iterative process with a number of steps bounded by the ambient dimension. This result, Theorem 7.1, is best stated in terms of nested normals, a concept that we introduce in Definition 6.3, and motivate throughout Section 6. For this reason, we defer it to Section 7.3. For the moment, we state a direct corollary which does not require this object.
@@
Now that the joint supporting subspace has been thoroughly identified, we move on to describing its core property. Recall that for two convex functions $f,g:\mathbb{R}^{n}\to\bar{\mathbb{R}}$, a common qualification condition is that for convex sets $C:=\operatorname{dom}(f),D:=\operatorname{dom}(g)$,
@@
It was shown in [rockafellarConvexAnalysis1970, Theorem 11.3] that this notion is equivalent to a lack of proper separation of the domains. We define the notions of separation, and proper separation below. We define these notions in affine subspaces, an intuitive generalization of separation in Euclidean spaces.
@@
Given an affine subspace $V\subseteq\mathbb{R}^{n}$, two convex sets $C,D\subseteq V$ are said to be separated in $V$ when there exists some $u\in(V-V)\setminus\{0\}$ such that
$\forall c\in C,d\in D,\langle u,c\rangle\leq\langle u,d\rangle.$ (4)
@@
For an affine subspace $V\subseteq\mathbb{R}^{n}$, convex sets $C,D\subseteq\mathbb{R}^{n}$ are properly separated in $V$ when there exists some $u\in(V-V)\setminus\{0\}$ such that Equation 4 holds, and additionally such that there exists some pair $(c,d)\in C\times D$ for which the inequality in Equation 4 is strict.
Proper separation excludes the case where the two sets are merely separated by virtue of being both contained in a single hyperplane. Since separation follows from proper separation, and lack of proper separation is equivalent to intersection of relative interiors (by [rockafellarConvexAnalysis1970, Theorem 11.3]), the contrapositive statement is that
$\text{lack of separation}\implies\text{intersection of relative interiors}.$ (5)
In the joint supporting subspace there is always a “lack of separation” of the convex sets.
Theorem 4.2 (Qualification conditions always hold in the joint supporting subspace).
For two convex sets $C,D\subseteq\mathbb{R}^{n}$ with $T_{a}:=T_{a}(C,D)$, $C\cap T_{a}$ and $D\cap T_{a}$ are not separated in $T_{a}$.
The proof of Theorem 4.2 is the subject of Section 7.1.
Corollary 4.5 (rint qualification holds in joint supporting subspace).
For any convex sets $C,D\subseteq\mathbb{R}^{n}$ such that $C\cap D\neq\emptyset$,
@@
Proof of Corollary 4.5.
The result follows from Theorem 4.2 by Equation 5, with the caveat that we defined separation in affine subspaces, which are not necessarily vector spaces, while in [rockafellarConvexAnalysis1970, theorem 11.3] this notion is defined in $\mathbb{R}^{n}$. Yet the argument still holds because the notions of separation, proper separation, and relative interiors are invariant under translation of $T_{a}$ to $T$. For any point $x_{0}\in T_{a}$, $C\cap T_{a}$ is (properly) separated from $C\cap T_{a}$ iff $C\cap T_{a}-x_{0}$ is (properly) separated from $D\cap T_{a}-x_{0}$ in $T$. A similar equivalence holds for intersection of relative interiors. ∎
@@ 5 Main results
With Corollary 4.5 we extend standard results of convex analysis by first localizing to $T_{a}$. We thus generalize the following beyond qualification conditions: subdifferential additivity; normal cones of the intersection of convex sets; attained infimal convolution for the conjugate of a sum $(f+g)^{*}$; the subdifferential chain rule; Fenchel-Rockafellar duality; and KKT conditions.
The next result implies that when standard qualification conditions hold, the classical formulation of our generalizations is trivially recovered.
@@
If the relative interiors of two convex sets $C,D\subseteq\mathbb{R}^{n}$ intersect, then
$C,D\subseteq T_{a}(C,D).$
5.1 Subdifferential calculus
Corollary 5.1 (Subdifferential sum rule).
For $f,g:\mathbb{R}^{n}\to\bar{\mathbb{R}}$ proper convex, $x\in\mathbb{R}^{n}$, and $T_{a}=T_{a}(\operatorname{dom}(f),\operatorname{dom}(g))$,
$\partial(f+g)(x)=\partial(f+\iota_{T_{a}})(x)+\partial(g+\iota_{T_{a}})(x).$
Proof of Corollary 5.1.
By Corollary 4.5 applied to $\operatorname{dom}(f)$ and $\operatorname{dom}(g)$, the result follows by application of [rockafellarConvexAnalysis1970, Theorem 23.8] to the functions $f+\iota_{T_{a}}$ and $g+\iota_{T_{a}}$. ∎
Corollary 5.2 (Normal cone of the intersection of convex sets).
For $C,D\subseteq\mathbb{R}^{n}$ convex, and $x\in C\cap D$, for $T_{a}=T_{a}(C,D)$,
@@ 5.1 Subdifferential calculus
Proof of Corollary 5.2.
With Corollary 4.5 we can use [rockafellarConvexAnalysis1970, Corollary 23.8.1], from which we find that
@@
In results that follow, we abuse notation by associating a matrix $A\in\mathbb{R}^{m\times n}$ with the corresponding linear map $A:\mathbb{R}^{n}\to\mathbb{R}^{m}$. For a set $S\subseteq\mathbb{R}^{n}$, we denote by $AS$ the set image $A(S)$, and by $A^{-1}(\cdot)$ the preimage of the linear map. We make no assumption on the invertibility of the matrix $A$.
Corollary 5.3 (Subdifferential chain rule).
Let $g\in\mathbb{R}^{m}\to\bar{\mathbb{R}}$ be a proper convex function, and $A\in\mathbb{R}^{m\times n}$ a matrix. Then for $T:=T(\operatorname{dom}(g),\operatorname{\mathrm{range}}(A))$,
$\partial(g\circ A)(x)=A^{T}\partial(g+\iota_{T})(Ax).$
Proof of Corollary 5.3.
@@
If $\operatorname{dom}(g)\cap\operatorname{\mathrm{range}}(A)=\emptyset,$ then the result holds trivially because $g(Ax)\equiv\infty$, and $(g+\iota_{T_{a}})(Ax)\equiv\infty$ as well. Otherwise, by Corollary 4.5, $\operatorname{relint}\operatorname{dom}(g+\iota_{T_{a}})\cap\operatorname{\mathrm{range}}(A)\neq\emptyset$, which lets us apply [rockafellarConvexAnalysis1970, Theorem 23.9] to the composition $g+\iota_{T_{a}}]\circ A$. Additionally, note that the only face of $\operatorname{\mathrm{range}}(A)$ is itself, whence by Lemma 4.1 $\operatorname{\mathrm{range}}(A)\subseteq T_{a}$. Since $0\in\operatorname{\mathrm{range}}(A)$, this in fact implies that $T_{a}=T$, and the result follows. ∎
5.2 Infimal convolution
$(f\,\square g)(z):=\inf_{x+y=z}f(x)+g(y).$
In the following results, we denote by $f^{*}$ the convex conjugate of $f$.
Corollary 5.4 (Attained infimal convolution).
For two proper convex functions $f,g:\mathbb{R}^{n}\to\bar{\mathbb{R}}$, and $T_{a}:=T_{a}(\operatorname{dom}(f),\operatorname{dom}(g))$,
$(f+g)^{*}=(f+\iota_{T_{a}})^{*}\square(g+\iota_{T_{a}})^{*},$
where the infimum in the infimal convolution is attained.
Proof of Corollary 5.4.
Whenever $\operatorname{dom}(f)\cap\operatorname{dom}(g)=\emptyset$, we have that $T_{a}=\emptyset$, in which case the equality holds trivially, with the infimum attained at any $y\in\mathbb{R}^{n}$. Otherwise, by Corollary 4.5 the result follows from [rockafellarConvexAnalysis1970, Theorem 20.1]. ∎
5.3 Convex optimization
We next examine the following general convex program.
Problem 5.1 (Convex program).
Given lower semicontinuous proper convex functions $f:\mathbb{R}^{n}\to\bar{\mathbb{R}}$, $g:\mathbb{R}^{m}\to\bar{\mathbb{R}}$, and a matrix $A\in\mathbb{R}^{m\times n}$,
$\mathop{\mathrm{minimize}}_{x\in\mathbb{R}^{n}}f(x)+g(Ax).$
A standard regularity condition for this problem is that
$A(\operatorname{relint}\operatorname{dom}(f))\cap\operatorname{relint}\operatorname{dom}(g)\neq\emptyset,$
which is a weakening of Slater’s condition. A program with this property is called strongly consistent [rockafellarConvexAnalysis1970, p. 300].
The next result shows the sense in which 5.1 can be localized to the joint facial subspace $T_{a}$, guaranteeing strong consistency whenever the program is feasible, while retaining all the crucial information of the original program.
Theorem 5.1 (Strong consistency of the localized program).
For $f:\mathbb{R}^{n}\to\bar{\mathbb{R}}$ and $g:\mathbb{R}^{m}\to\bar{\mathbb{R}}$ lower semicontinuous proper convex functions, a matrix $A\in\mathbb{R}^{n\times m}$, $T_{a}:=T_{a}(A\operatorname{dom}(f),\operatorname{dom}(g))$, and any $x\in\mathbb{R}^{n}$,
$f(x)+g(Ax)=(f(x)+\iota_{A^{-1}(T_{a})}(x))+(g(x)+\iota_{T_{a}}(x)).$ (6)
And whenever $\operatorname{dom}(f+g\circ A)\neq\emptyset$,
$A\operatorname{relint}\operatorname{dom}(f+\iota_{A^{-1}(T_{a})})\cap\operatorname{relint}\operatorname{dom}(g+\iota_{T_{a}})\neq\emptyset.$